RC Meeting, Tuesday, August 31, 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meeting at 10:00 am, (remote)

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Accelerator Status

  • Maintenance Day Today, followed by first beam to Hall B
  • Box supplies 1-7 ready. 8-10 being worked on
  • 5-pass setup September 2nd - 5th
  • Tune Beam to C September 6th
  • Expected start of Physics: September 7th

Hall Status

  • Lead Collimator installed
  • Target pumped down. Possible cool down Wednesday. Need to hear back from Cryo. Training can start when target is cooled.
  • A/C work being performed in the Hall. After lift will be taken out and Hall can be locked up. Walter might do his checklist Wednesday
  • Steve L. asked Walter to pump down on a sweaty cryo line
  • Limit switch for SHMS needs to be moved to 5.49 degree position (today?)
  • Steve was asked to put the spectrometers at starting position. SHMS=7.5deg, HMS=12.5deg

Notices in Effect for Shift Crew

  • Please read safety documentation (https://hallcweb.jlab.org/safety-docs/Fall2021/) and sign in the "yellow binder" in the Counting House
  • In the near future....vaccination cards must be submitted to Occupational Medicine in order to get site access. Or unvaccinated and undeclared users will need to have a negative test with 3 days of arrival to gain access and will need to be tested weekly. Occupation Medicine will be informing the user community as soon as the details are finalized.

Current Issues

  • Missing FADC information in PMT#2 of the noble gas cerenkov. Signals seen on scope. Investigation ongoing
  • Interlock tripping on HMS Q1 [https://logbooks.jlab.org/entry/3896458} Steve took a look yesterday. Mark will exercise magnets today.
  • Target motion camera should be installed. Waiting to hear back from Joe B.
  • Checklist need to be done to lock up the Hall (Brad, Walter, Dave G, Steve L). Will wait until after tomorrows scheduling meeting to schedule.
  • Shifts need to be filled for next week (Monday SWING) [1]
  • Missing links in wiki run plan [2]
  • Checklist not posting to HCLOG [3]